The nucleus of 8Be, which consists of 4 protons and 4 neutrons, is very unstable and spontaneously breaks into two alpha particles (helium nuclei, each consisting of 2 protons and 2 neutrons). (a) What is the force between the two alpha particles when they are 4.46 10-15 m apart? 46.33111464N (b) What is the initial magnitude of the acceleration of the alpha particles due to this force? Note that the mass of an alpha particle is 4.0026 u Enter a number differs significantly from the correct answer. Rework your solution from the beginning and check each step carefully. m/s2 Need Help?ReadIt Submit Answer Save ProgressPractice Another Version 0/1 points I Previous Answers SerCP11 15.2.P006. My Notes Ask Your Teacher A molecule of DNA (deoxyribonucleic acid) is 2.00 (um long. The ends of the molecule become singly ionized: negative on one end, positive on the other. The helical molecule acts like a spring and compresses 1.21% upon becoming charged. Determine the effective spring constant of the molecule. Your response is within 10% of the correct value. This may be due to roundoff error, or you could have a mistake in your calculation. Carry out all intermediate results to at least four-digit accuracy to minimize roundoff error. N/m Need Help?Read It

Please give me the answer as whole number not as power to the ten. (ex 10^23 or 10^-23).
Answer #1

1b] magnitude of acceleration a = F/m = 46.3311/[4.0026*1.66054e-27]

= 6.9707752e27 m/s^2 = 6.97*10^27 m/s^2

2] The net force on one end : Spring force - colombs force = 0

kx = Kq^2/d^2

k*[2e-6*1.21e-2] = 9e9*1.6e-19^2/[2e-6 - 2e-6*1.21e-2]^2

k = 2.44*10^-9 N/m answer